# Emergent topological re-entrant phase transition in a generalized quasiperiodic modulated Su-Schrieffer-Heeger model

### v1

@@ I Introduction
Topological insulators (TI), as a system with unique transport properties, constitute one of the most important research directions in condensed matter physics and quantum computation x1 ; x2 ; x3 ; x4 ; x5 ; x6 ; x7 ; x8 ; x10 ; x13 ; x14 ; x15 ; x16 ; x17 ; x18 ; x19 ; x20 ; 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 . Previous studies have shown that the non-trivial edge states in topological systems feature good robustness, which ensures that the edge current can still maintain its original state in the case of weak disorder or defects. However, when disorder becomes very strong, the topological system will undergo Anderson phase transition 03 . In other words, in the case of strong disorder, the topological edge current will be destroyed to make the topological system a gapless Anderson insulator, where the corresponding bulk states show the characteristics of localized states. In addition, recent studies have revealed that moderate disorder can achieve a transition from a trivial phase to a non-trivial phase, and this topological system induced by disorder is called topological Anderson insulator (TAI) 001 ; traceb0 ; 04 ; 05 ; 06 ; 07 ; f6 ; f7 ; f9 ; f10 ; h01 ; h02 ; h03 ; h04 ; h05 ; h06 ; h07 ; h08 ; h1 .
In recent years, TAIs and its related fields have been greatly developed, and a series of milestone achievements have been scored. Theoretically, in addition to the standard TAI, the following systems have also been predicted: $\mathbb{Z}_{2}$ topological Anderson insulators e1 , Topological Anderson amorphous insulator e2 , Higher-order Topological Anderson Insulators e3 ; h2 , Topological inverse Anderson insulator e4 , etc e5 ; e6 . Besides, with the increasing popularity of non-Hermitian research nh00 ; nh001 ; nh1 ; nh002 ; nh01 ; nh02 ; nh03 , the study on TAI has gradually extended to non-Hermitian systems z1 ; z02 ; z03 ; z04 . Experimentally, TAI has been realized in a variety of artificial systems, including ultra-cold atoms z2 ; z3 ; zz0 , photonic/phononic system z4 ; z5 ; z6 , superconducting system 002 , and electric circuits z7 , etc.
Re-entrant phase transition (REPT), on the other hand, refers to the process in which the system starts from a phase and returns to the same phase by monotonically manipulating a certain parameter z8 . Recently, the REPT of localized phase has been discovered in Aubry-André model a0 ; a1 . Moreover, topological REPT has also been reported recently tr1 . So far, although there are many researches on TI and TAI, few work has been done on topological REPT in TI and TAI tr2 . This paper is devoted to the study of REPT phenomena in TI and TAI system.
@@ II MODEL and Key quantities
[…]
@@
where $t_{1}$ is the intracell hopping strength, $\lambda$ is the strength of quasiperiodic modulation, $b$ is the structure factor, which is the key parameter to control the quasiperiodic modulation bounded ($b<1$) or unbounded ($b\geq 1$). $\theta$ is an additional phase shift, and $\alpha$ is an irrational number. When $\lambda=0$, the Hamiltonian Eq. (1) reduces to the standard SSH model ssh1 . The system represents a trivial (non-trivial) topology under the condition of $t_{1}<t_{2}$ $(t_{1}>t_{2})$. Without loss of generality, we set $t_{2}=1$ as the energy unit. We take $\alpha=(\sqrt{5}-1)/2$ and $\theta=0$. In the numerical calculation, we choose the system size $N=L/2=610$ with $L$ being the total lattice number. Such size is large enough for self-averaging and one can safely ignore the finite size effect (see Appendix A for details).
@@
where $\text{Q}={\textstyle\sum_{j=1}^{N}}(\left|j\right\rangle\left\langle j\right|-\left|\tilde{j}\right\rangle\left\langle\tilde{j}\right|)$ is the corresponding open-boundary matrix and $\left|\tilde{j}\right\rangle=\Gamma^{-1}\left|j\right\rangle$. One can directly obtain the matrix Q by solving the eigenequations $H\left|j\right\rangle=E_{j}\left|j\right\rangle$, where $E_{j}$ and $\left|j\right\rangle$ correspond to the eigenenergies and eigenstates, respectively. $\Gamma=I_{N}\otimes\sigma_{z}$ denotes the chiral symmetry operator with the identity matrix $I_{N}$ and the Pauli matrix $\sigma_{z}$. The key operator
@@
is the coordinate operator. The symbol $\operatorname{Tr}^{{}^{\prime}}$ means the trace over the middle interval of the full lattice with the length $L^{{}^{\prime}}=L/2=N$ (see Appendix A for details). For example, under the condition of the primitive cells’ number $N=6$. Since a primitive cell contains two sublattices, the total number of sites is $12$. The corresponding matrix in Eq. (4) is a $12\times 12$ matrix. Then, $\operatorname{Tr}^{\prime}$ represents the trace of a $6\times 6$ submatrix formed by selecting the middle region of the original matrix (the region of rows $4-9$$\times$ column $4-9$).
@@ III Topological REPT for the BOUNDED case
Let’s start with the bounded case, i.e., the case where $b<1$. Without loss of generality, we fix $b=0.9$ to show how the topological properties of the system change with $\lambda$ and $t_{1}$. The corresponding results are shown in Fig. 2(a). We find that, unlike traditional topological phase transitions, topological REPT phenomena will occur in the range of parameter $t_{1}\in[0.7,1]$ for the model of Eq. (1). Taking $t_{1}=0.8$ (red dashed line) as an example. One can see that topological REPT phenomenon will emerge as the quasiperiodic strength $\lambda$ increases. Specifically, the system starts from a traditional TI phase. With the increase of $\lambda$, the system first enters the trivial phase, and then enters the TAI phase. Finally, when the disorder strength completely dominates, the system will enter the trivial region due to Anderson localization. We refer to this phenomenon of re-entrant from traditional TI into TAI as the type-I of topological REPT.
@@
Figure 3: Density distribution of the $N$-th and $N+1$-th eigenstates under the condition of $\lambda=0,~{}0.5,~{}2,~{}2.5$ (marked by green squares in Fig. 2). Throughout, we set $t_{1}=0.8$.
@@ IV TOPOLOGICAL REPT FOR THE UNBOUNDED CASE
Figure 5: Density distribution of the $N$-th and $N+1$-th eigenstates under the condition of $\lambda=0,~{}0.5,~{}1.5,~{}3,~{}3.5$ (marked by green squares in Fig. 4). Throughout, we set $t_{1}=1.2$.
V CONCLUSIONS
@@ Appendix A FINITE SIZE EFFECT
With a fixed primitive size $N=610$, we show the case of trace with the whole matrix ($L$) and the half matrix blocks ($L/2$), respectively. It is not difficult to find that the winding number is best calculated using the trace of the central $L/2$ block. However, when we choose the trace of the whole matrix to calculate the winding number, there will be obvious errors in the result. This is because as a topology marker, the edge elements of the diagonal matrix in Eq. (9) lead to inaccurate results [see Fig. 7(b)]. Therefore, as in previous studies  001 ; traceb0 , in main text, we also select the central $L/2$ region to calculate the winding number.

### v2

@@ I Introduction
Topological insulators (TI), as a system with unique transport properties, constitute one of the most important research directions in condensed matter physics and quantum computation x1 ; x2 ; x3 ; x4 ; x5 ; x6 ; x7 ; x8 ; x10 ; x13 ; x14 ; x15 ; x16 ; x17 ; x18 ; x19 ; x20 ; 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 . Previous studies have shown that the non-trivial edge states in topological systems feature good robustness, which ensures that the edge current can still maintain its original state in the case of weak disorder or defects. However, when disorder becomes very strong, the topological system will undergo Anderson phase transition 03 . In other words, in the case of strong disorder, the topological edge current will be destroyed to make the topological system a gapless Anderson insulator, where the corresponding bulk states show the characteristics of localized states. In addition, recent studies have revealed that moderate disorder can achieve a transition from a trivial phase to a non-trivial phase, and this topological system induced by disorder is called topological Anderson insulator (TAI) 001 ; traceb0 ; 04 ; 05 ; 06 ; 07 ; f6 ; f7 ; f9 ; f10 ; h01 ; h02 ; h03 ; h04 ; h05 ; h06 ; h07 ; h08 ; h1 ; h09 .
In recent years, TAIs and its related fields have been greatly developed, and a series of milestone achievements have been scored. Theoretically, in addition to the standard TAI, the following systems have also been predicted: $\mathbb{Z}_{2}$ topological Anderson insulators e1 , Topological Anderson amorphous insulator e2 , Higher-order Topological Anderson Insulators e3 ; h2 , Topological inverse Anderson insulator e4 , etc e5 ; e6 ; e7 . Besides, with the increasing popularity of non-Hermitian research nh00 ; nh001 ; nh1 ; nh002 ; nh01 ; nh02 ; nh03 ; nh04 ; nh05 , the study on TAI has gradually extended to non-Hermitian systems z1 ; z02 ; z03 ; z04 . Experimentally, TAI has been realized in a variety of artificial systems, including ultra-cold atoms z2 ; z3 ; zz0 , photonic/phononic system z4 ; z5 ; z6 , superconducting system 002 , and electric circuits z7 , etc.
Re-entrant phase transition (REPT), on the other hand, refers to the process in which the system starts from a phase and returns to the same phase by monotonically manipulating a certain parameter z8 . Recently, the REPT of localized phase has been discovered in Aubry-André model a0 ; a1 and SSH model 519 ; 512 , etc 520 ; 521 . Moreover, topological REPT has also been reported recently tr1 . So far, although there are many researches on TI and TAI, few work has been done on topological REPT in TI and TAI tr2 . This paper is devoted to the study of REPT phenomena in TI and TAI system.
@@ II MODEL and Key quantities
$H=\sum_{n=1}^{N}\left(t_{1}^{{}^{\prime}}a_{n}^{\dagger}b_{n}+t_{2}a_{n+1}^{\dagger}b_{n}+\text{ H.c. }\right),$ (1)
@@
where $t_{1}$ is the intracell hopping strength, $\lambda$ is the strength of quasiperiodic modulation, $b$ is the structure factor, which is the key parameter to control the quasiperiodic modulation bounded ($\left|b\right|<1$) or unbounded ($\left|b\right|\geq 1$). $\theta$ is an additional phase shift, and $\alpha$ is an irrational number. When $\lambda=0$, the Hamiltonian Eq. (1) reduces to the standard SSH model ssh1 . The system represents a trivial (non-trivial) topology under the condition of $t_{1}<t_{2}$ $(t_{1}>t_{2})$. Without loss of generality, we set $t_{2}=1$ as the energy unit. We take $\alpha=(\sqrt{5}-1)/2$ and $\theta=0$. In the numerical calculation, we choose the system size $N=L/2=610$ with $L$ being the total lattice number. Such size is large enough for self-averaging and one can safely ignore the finite size effect (see Appendix A for details).
@@
[…]
@@
is the coordinate operator. The symbol $\operatorname{Tr}^{{}^{\prime}}$ means the trace over the middle interval of the full lattice with the length $L^{{}^{\prime}}=L/2=N$ (see Appendix A for details). For example, under the condition of the primitive cells’ number $N=6$. Since a primitive cell contains two sublattices, the total number of sites is $12$. The corresponding matrix in Eq. (4) is a $12\times 12$ matrix. Then, $\operatorname{Tr}^{\prime}$ represents the trace of a $6\times 6$ submatrix formed by selecting the middle region of the original matrix (the region of rows 494-9 $\times$ column 494-9).
@@ III Topological REPT for the BOUNDED case
Let’s start with the bounded case, i.e., the case where $\left|b\right|<1$. Without loss of generality, we fix $b=0.9$ to show how the topological properties of the system change with $\lambda$ and $t_{1}$. The corresponding results are shown in Fig. 2(a). We find that, unlike traditional topological phase transitions, topological REPT phenomena will occur in the range of parameter $t_{1}\in[0.7,1]$ for the model of Eq. (1). Taking $t_{1}=0.8$ (red dashed line) as an example. One can see that topological REPT phenomenon will emerge as the quasiperiodic strength $\lambda$ increases. Specifically, the system starts from a traditional TI phase. With the increase of $\lambda$, the system first enters the trivial phase, and then enters the TAI phase. Finally, when the disorder strength completely dominates, the system will enter the trivial region due to Anderson localization. We refer to this phenomenon of re-entrant from traditional TI into TAI as the type-I of topological REPT.
@@
Figure 3: Density distribution of the NN-th and $N+1$-th eigenstates under the condition of $\lambda=0,~0.5,~2,~2.5$ (marked by green squares in Fig. 2). Throughout, we set $t_{1}=0.8$.
@@ IV TOPOLOGICAL REPT FOR THE UNBOUNDED CASE
Figure 5: Density distribution of the NN-th and $N+1$-th eigenstates under the condition of $\lambda=0,~0.5,~1.5,~3,~3.5$ (marked by green squares in Fig. 4). Throughout, we set $t_{1}=1.2$.
V Discussions and CONCLUSIONS
Topological REPT has recently received much attention because of its important theoretical and potential application value. Re-entrant integer quantum Hall effect, re-entrant quantum spin Hall effect, re-entrant quantum anomalous Hall effect, and re-entrant topological superconductivity have been successively discovered z9 ; re51 ; z10 ; re52 ; 513 ; 514 ; zp11 ; zp12 ; zp13 ; 516 . The systems involved are also diverse, including traditional topological systems z9 ; re51 ; z10 ; re52 ; 513 ; 514 , topological superconductors zp11 ; zp12 ; zp13 ; 516 , hybrid bridging systems zp14 , non-Hermitian systems 515 , and so on 517 ; pl1 ; 518 . However, there are two differences between the topological re-entrant seen in these systems and the REPT phenomenon proposed in this paper, i.e., the inducers of REPT as well as the topological phases involved in REPT. For example, the reentrant integer quantum Hall effect occurs in a two-dimensional degenerate electron gas as the magnetic field changes z9 ; re51 ; z10 ; re52 , and the bridging model of Kitaev chain and Haldane chain realize REPT from Majorana zero mode phase to Z2 topologically ordered phase through pairing interaction zp14 . It can be seen clearly that the former is induced by magnetic fields, and the latter is induced by pairing interactions. While, both are fundamentally different from the phenomenon studied in this paper that is controlled by disorder. In addition, most of the previous studies involved traditional topological phases, while this work focuses on the REPT of the topological (Anderson) insulator phases in a weakly disordered region, and this non-traditional object of study constitutes the other major difference.
@@ Appendix A FINITE SIZE EFFECT
With a fixed primitive size $N=610$, we show the case of trace with the whole matrix ($L$) and the half matrix blocks ($L/2$), respectively. It is not difficult to find that the winding number is best calculated using the trace of the central $L/2$ block. However, when we choose the trace of the whole matrix to calculate the winding number, there will be obvious errors in the result. This is because as a topology marker, the edge elements of the diagonal matrix in Eq. (9) lead to inaccurate results [see Fig. 7(b)]. Therefore, as in previous studies  001 ; traceb0 ; Fi1 , in main text, we also select the central $L/2$ region to calculate the winding number.
@@ Appendix B TOPOLOGICAL PHASE DIAGRAM FOR THE BOUNDED TO UNBOUNDED CASE
Appendix C FIDELITY SUSCEPTIBILITY
The fidelity susceptibility $\chi_{F}$ reads FS1
$\chi_{F}=-\frac{\partial^{2}F}{\partial(\delta\lambda)^{2}}=\lim_{\delta\lambda\rightarrow 0}\frac{-2\log F}{(\delta\lambda)^{2}},$ (11)
with
$F\left(\Psi(\lambda),\Psi\left(\lambda^{\prime}\right)\right)=\left|\operatorname{det}\left(\boldsymbol{\Phi}_{\lambda}^{\dagger}\boldsymbol{\Phi}_{\lambda^{\prime}}\right)\right|,$ (12)
where $\Phi_{\lambda}$ is a $N\times M$ matrix containing the M filled single-particle eigenstates in its columns. We show how the fidelity changes with $\lambda$ for three different sizes. It is easy to find that the peak of the fidelity at critical points becomes sharper and sharper with the increase of the size, which indicates the appearance of ciritical points (see Fig. 9).
Figure 9: Fidelity susceptibility $\chi_{F}$ versus $\lambda$ with $b=0.9$ and $t_{1}=0.8$ (a), and $b=1.5$ and $t_{1}=1.2$ (b). System sizes are marked.